OpenStudy (anonymous):

A natural bridge in Utah :D has a span of 60 feet more than than its height were 200 feet less, it would be 1/5 of its span. Find the height and the span of the bridge. Show all work

OpenStudy (anonymous):

problem not clear. repost it

OpenStudy (dumbcow):

i think what you are saying is given h for height and s for span: s = h+60 s/5 = h -200 in that case solve system using substitution (h+60)/5 = h-200 h+60 = 5h -1000 -4h = -1060 h = 265 s = 265 + 60 = 325